### v1

# Quantum Byzantine Multiple Access Channels

Abstract
In communication theory, attacks like eavesdropping or jamming are typically assumed to occur at the channel level, while communication parties are expected to follow established protocols. But what happens if one of the parties turns malicious? In this work, we investigate a compelling scenario: a multiple-access channel with two transmitters and one receiver, where one transmitter deviates from the protocol and acts dishonestly. To address this challenge, we introduce the Byzantine multiple-access classical-quantum channel and derive an achievable communication rate for this adversarial setting.
I Introduction
The rapid advancements in modern communication systems open up new possibilities while simultaneously exposing us to numerous cybersecurity threats. Many of these systems remain imperfect and vulnerable to hostile attacks, necessitating robust protection against malicious interference.
A prominent example is the forthcoming 6G communication standard, which will integrate key technological infrastructure for the metaverse and other digital-twin applications. Given its profound impact on future applications, particularly in sensitive aspects of human well-being, ensuring a high degree of trustworthiness is imperative ([10, 16, 17, 18]).
Communication models that account for a jammer attempting to disrupt proper communication between legitimate parties-often represented by the arbitrarily varying channel-have been extensively studied in both classical ([9], [2]) and quantum information theory ([5], [7], [4]).
A multiple-access channel is a communication channel where two or more senders transmit information to a common receiver. The optimal coding strategies and capacity regions for classical multiple-access channels have been established in [3] and [26], while those for multiple-access quantum channels have been derived in [34] and [35].
A multiple-access channel with a jammer of known identity is referred to as an arbitrarily varying multiple-access channel. The capacity regions of classical arbitrarily varying multiple-access channels were determined in [24]. The symmetrizability conditions for these channels were studied in [19] and [6]. In the classical-quantum setting, [14] established the capacity region by introducing a novel coding technique that does not rely on conditional typical subspace projectors.
However, modern digital communication systems are increasingly vulnerable to cyber threats, not only from external attackers such as hackers or government agencies but also from within. In some cases, protection against external threats alone is insufficient, as certain parties may act dishonestly or even attempt to disrupt the communication of honest participants.
Unlike arbitrarily varying channel models, the Byzantine model considers scenarios where one of the legitimate users may behave adversarially. Byzantine attacks on network coding have been extensively studied in [25], [23], and [32].
Figure 1: classical-quantum 3-user byzantine-multiple access channel, one user may be adversarial.
In this work, we explore the quantum counterpart of a recently introduced communication model: the Byzantine multiple-access channel, first studied in [29] and [30]. Specifically, we investigate a classical-quantum multiple-access channel where one of the transmitters may act adversarially. Instead of transmitting legitimate messages, the adversarial sender may inject malicious inputs as a jamming attack to disrupt communication.
A key challenge in this setting is that the receiver does not know which sender is adversarial. Despite this uncertainty, we require that the messages of honest senders be reliably decoded. The capacity region of the classical Byzantine multiple-access channel was established in [29], while [30] introduced a technique for identifying the adversarial sender.
When extending communication models to the quantum domain, several new challenges arise. In quantum systems, every measurement can alter the quantum state, meaning that disruptions may not only come from an adversarial sender’s inputs but also from previously applied decoding operations (cf. Section III-A).
Moreover, while the capacity of a classical arbitrarily varying channel can be achieved using a deterministic code-without relying on any randomness (cf. [13] and [6], where a deterministic code for an arbitrarily varying multiple-access channel was constructed)- the same does not hold for quantum arbitrarily varying channels, where only random coding techniques have been developed so far (cf. [5] and [14]).
As a result, many of the sophisticated techniques from [29] cannot be directly extended to quantum Byzantine multiple-access channels. Similarly, the identification technique proposed in [30] is not applicable: while a sequence of classical symbols can be easily verified against a set of expected legal channel outputs, a “false” quantum state may still have a nonzero probability of being accepted by the decoding operators for legitimate messages. To overcome these challenges, we must develop a novel coding approach.
II Preliminaries
Let $\rho_{1}$ and $\rho_{2}$ be Hermitian operators on a finite-dimensional complex Hilbert space $A$. We write $\rho_{1}\geq\rho_{2}$ (or equivalently, $\rho_{2}\leq\rho_{1}$) if $\rho_{1}-\rho_{2}$ is positive semidefinite. For a finite-dimensional complex Hilbert space $A$, the set of density operators on $A$ is denoted by
$\mathcal{S}(A):=\left\{\rho\in\mathcal{L}(A)\mid\rho\text{ is Hermitian, }\rho\geq 0,\text{ and }\mathrm{tr}(\rho)=1\right\},$
where $\mathcal{L}(A)$ is the set of linear operators on $A$, and $0_{A}$ denotes the null matrix on $A$. Note that every operator in $\mathcal{S}(A)$ is bounded.
Let $X$ be a finite alphabet, and $A$ be a finite-dimensional complex Hilbert space. A classical-quantum channel is a map $\mathtt{W}:X\rightarrow\mathcal{S}(A)$, where $x\mapsto\mathtt{W}(x)$.
For a quantum state $\rho\in\mathcal{S}(A)$, the von Neumann entropy of $\rho$ is denoted by $S(\rho)=-\mathrm{tr}(\rho\log\rho)$.
Given a set $X$, a Hilbert space $A$, a classical-quantum channel $\mathtt{V}:X\rightarrow\mathcal{S}(A)$, and a probability distribution $P$ on $X$, the conditional von Neumann entropy is defined as
$S(\mathtt{V}|P):=\sum_{x\in X}P(x)S\left(\mathtt{V}(x)\right).$
The quantum relative entropy between $\rho$ and $\sigma$ is defined as
$D(\rho\parallel\sigma):=\mathrm{tr}\left(\rho\left(\log\rho-\log\sigma\right)\right),$
if $\text{supp}(\rho)\subset\text{supp}(\sigma)$, and $D(\rho\parallel\sigma)=\infty$ otherwise.
For a set $X$, a Hilbert space $A$, a classical-quantum channel $\mathtt{V}:X\rightarrow\mathcal{S}(A)$, and a probability distribution $P$ on $X$, the quantum mutual information of the channel $\mathtt{V}$ with input distribution $P$ is defined as
$I(P;\mathtt{V}):=\sum_{x\in X}P(x)D\left(\mathtt{V}(x)\parallel\sum_{x^{\prime}\in X}P(x^{\prime})\mathtt{V}(x^{\prime})\right).$
Let $\Phi:=\{\rho_{x}:x\in X\}$ be a set of quantum states labeled by elements of $X$. For a probability distribution $P$ on $X$, the Holevo quantity, or Holevo information, is defined as
$\chi(P;\Phi):=S\left(\sum_{x\in X}P(x)\rho_{x}\right)-\sum_{x\in X}P(x)S(\rho_{x}).$
For a set $X$ and a Hilbert space $A$, let $\mathtt{V}:X\rightarrow\mathcal{S}(A)$ be a classical-quantum channel. For a probability distribution $P$ on $X$, the Holevo quantity of the channel $\mathtt{V}$ with input distribution $P$ is defined as
$\chi(P;\mathtt{V}):=S\left(\sum_{x\in X}P(x)\mathtt{V}(x)\right)-S(\mathtt{V}|P).$
Notice that when we choose an “optimal” set $\Phi:=\{\rho_{x}:x\in X\}$ (in the sense that it maximizes the quantum mutual information and the Holevo quantity), we have (cf. [21])
$I(P;\mathtt{V})=\chi(P;\mathtt{V})\quad\text{.}$
Definition II.1.
A set of classical-quantum channels $\{\mathbf{W}(\cdot,t):t\in\Theta\}$ is called an arbitrarily varying classical-quantum channel when the state $t$ varies from symbol to symbol in an arbitrary manner.
We say that the arbitrarily varying classical-quantum channel $\{\mathbf{W}(\cdot,t):t\in\Theta\}$ is symmetrizable if there exists a parametric set of distributions $\{\tau(\cdot\mid x):x\in X\}$ on $\Theta$ such that for all $x,x^{\prime}\in X$, the following condition holds:
$\sum_{t\in\Theta}\tau(t\mid x)\mathbf{W}(x^{\prime},t)=\sum_{t\in\Theta}\tau(t\mid x^{\prime})\mathbf{W}(x,t).$
Definition II.2.
Consider $k$ independent senders, with sender $i$ using an alphabet $X_{i}$ and an a priori probability distribution $P_{i}$. Let $A$ be a finite-dimensional complex Hilbert space.
A $k$-user Byzantine multiple-access classical-quantum channel $\mathtt{W}$ is a map
$\mathtt{W}:X_{1}\times\cdots\times X_{k}\rightarrow\mathcal{S}(A),$
where one of the senders may be adversarial, but the decoder does not know a priori which sender is adversarial.
Definition II.3.
A random code for a $k$-user Byzantine multiple-access classical-quantum channel $\mathtt{W}$ consists of an encoder
$\Bigl{\{}\{x_{i}^{n}(m_{i})^{\gamma_{i}}:\gamma_{i}\in\Lambda_{i}\}:m_{i}\in M_{i}\Bigr{\}}\subset X_{i}^{n}$
for every sender $i\in\{1,\dots,k\}$, and a set of positive operator-valued measure (POVM) decoder operators
$\Bigl{\{}\left\{D_{m_{i}}^{\gamma_{i}}:m_{i}\in M_{i}\right\}:\gamma_{i}\in\Lambda_{i}\Bigr{\}}$
for each sender’s message. These are parameterized by a random variable on a finite set $\Lambda_{i}$. The interpretation of this definition is that both sender $i$ and the receiver have access to the outcome of a random experiment uniformly distributed on $\Lambda_{i}$.
The average probability of decoding error for the message of sender $i$ is defined as
$1-\frac{1}{J_{n}}\frac{1}{|\Lambda_{i}|}\sum_{m_{i}\in M_{i}}\sum_{\gamma_{i}\in\Lambda_{i}}\mathrm{tr}\left({\mathtt{W}}^{\otimes n}\left(x_{i}^{n}(m_{i})^{\gamma_{i}}\right)D_{m_{i}}^{\gamma_{i}}\right).$
Definition II.4.
A non-negative number $R_{i}$ is an achievable rate for the message of sender $i$ in a Byzantine multiple-access classical-quantum channel $\mathtt{W}$ under random coding if, for every $\delta>0$, $\epsilon>0$, and sufficiently large $n$, there exists a random code such that
$\frac{\log|M_{i}|}{n}>R_{i}-\delta,$
and the average probability of decoding error for the message of sender $i$ is less than $\epsilon$ for any input sequence from the adversarial sender.
The closure of all achievable random rates for $\mathtt{W}$ is called the random capacity region of $\mathtt{W}$.
III Message Transmission over Quantum Byzantine Multiple Access Channel
III-A Necessity of Decoding Order
In the classical multiple-access channel scenario, it is evident that the order in which the messages of the senders are decoded does not matter. In fact, most classical multiple-access codes decode all messages simultaneously (cf. [3] and [26]). However, in the quantum multiple-access channel scenario, the previously decoded operator can affect the subsequent decoded messages, as every measurement may alter the channel outputs.
The standard solution to this issue in the quantum multiple-access channel scenario is based on the following observation. Suppose the first decoding operator can decode the first message with a probability sufficiently close to 1, and assume this decoding operator is constructed using subspace projections. In this case, we can apply the Tender Operator Lemma (cf. [33] and [28]) to show that this decoding operator does not significantly alter the channel outputs.
However, in the Byzantine multiple-access classical-quantum channel scenario, we cannot apply the Tender Operator trick, since we cannot guarantee that the first decoding operator will decode the first message with a sufficiently high probability, as the first sender may indeed be the adversarial one. Therefore, in this scenario, it is possible that if the first sender is adversarial, the first decoding operator may alter the channel outputs in such a way that the second message becomes undecodable. The following simple example illustrates this point:
We define the input alphabets as $X_{1}:=\{0,1\}$ and $X_{2}:=\{0,1,2\}$. Let $A\otimes B$ be a six-dimensional Hilbert space spanned by the orthonormal basis:
$\{|i\rangle^{A}\otimes|j\rangle^{B}:i\in\{0,1\},j\in\{0,1,2\}\}.$
Consider a two-user Byzantine multiple access classical-quantum channel $\mathtt{W}:X_{1}\times X_{2}\to\mathcal{S}(A\otimes B)$, defined by
$\mathtt{W}((i,j))=(|i\rangle^{A}\otimes|j\rangle^{B})(\langle j|^{B}\otimes\langle i|^{A}),\quad\forall i\in\{0,1\},j\in\{0,1,2\}.$
Suppose sender 1 transmits messages from the set $\{m_{0}^{1},m_{1}^{1}\}$, and sender 2 transmits messages from the set $\{m_{0}^{2},m_{1}^{2}\}$. That is, each sender’s message set consists of two possible messages.
Encoding Strategy
Each sender encodes their message as follows: For $k\in\{1,2\}$, sender $k$ simply transmits $i\in\{0,1\}$ as their channel input to represent message $m_{i}^{k}$.
Decoding Strategy
The receiver employs the following POVM to decode sender 1’s message:
$D_{0}^{(1)}:=\sum_{j=0}^{2}(|0\rangle^{A}\otimes|j\rangle^{B})(\langle j|^{B}\otimes\langle 0|^{A}),$
$D_{1}^{(1)}:=\sum_{j=0}^{2}(|1\rangle^{A}\otimes|j\rangle^{B})(\langle j|^{B}\otimes\langle 1|^{A}).$
In other words, the receiver measures only the $A$-part of the channel output, leaving the $B$-part unchanged.
The receiver employs the following POVM to decode the message of sender 2:
$D_{0}^{(2)}$ $=(|0\rangle^{A}\otimes|0\rangle^{B})(\langle 0|^{B}\otimes\langle 0|^{A})+(|1\rangle^{A}\otimes|0\rangle^{B})(\langle 0|^{B}\otimes\langle 1|^{A})$
$\quad+\frac{1}{2}(|0\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 0|^{A})+\frac{1}{2}(|1\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 1|^{A})$
$\quad+\frac{1}{2}(|0\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 1|^{A})+\frac{1}{2}(|1\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 0|^{A}),$
$D_{1}^{(2)}$ $=(|0\rangle^{A}\otimes|1\rangle^{B})(\langle 1|^{B}\otimes\langle 0|^{A})+(|1\rangle^{A}\otimes|1\rangle^{B})(\langle 1|^{B}\otimes\langle 1|^{A})$
$\quad+\frac{1}{2}(|0\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 0|^{A})+\frac{1}{2}(|1\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 1|^{A})$
$\quad-\frac{1}{2}(|0\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 1|^{A})-\frac{1}{2}(|1\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 0|^{A}).$
The receiver applies these operators to decode the message of sender 2.
Figure 2: Sender $1$ is trustworthy, while sender $2$ is adversarial. The receiver first decodes the message from sender $1$, followed by the message from sender $2$.
Figure 3: Sender $2$ is trustworthy, while sender $1$ is adversarial. The receiver first decodes the message from sender $1$, followed by the message from sender $2$.
First, assume that the receiver decides to decode the message of sender $1$ first. We distinguish three different cases:
Case 1a: Sender $1$ is trustworthy, while sender $2$ is adversarial. In this case, the receiver can decode sender $1$’s message with probability $1$, regardless of the input from sender $2$.
Case 1b: Both senders are trustworthy. The receiver first decodes sender $1$’s message with probability $1$. Since this decoding does not alter part $B$ of the channel outputs, the receiver can then decode sender $2$’s message with probability $1$ as well.
Case 1c: Sender $1$ is adversarial, while sender $2$ is trustworthy. Since the measurement operators $D_{0}^{(1)}$ and $D_{1}^{(1)}$ do not affect part $B$ of the channel outputs, the receiver can still decode sender $2$’s message with probability $1$, regardless of the input from sender $1$.
Thus, in this decoding order, we obtain a reliable code for the Byzantine multiple-access channel.
Figure 4: Sender $2$ is trustworthy, while sender $1$ is adversarial. The receiver first decodes sender $2$’s message, followed by sender $1$’s message.
Figure 5: Sender $1$ is trustworthy, while sender $2$ is adversarial. The receiver first decodes sender $2$’s message, followed by sender $1$’s message. In this case, the decoding operator for sender $2$’s message interferes with the decoding of sender $1$’s message.
Now, let us assume that the receiver decides to decode the message of sender $2$ first. Once again, we distinguish three different cases:
Case 2a: Sender $2$ is trustworthy, but sender $1$ is adversarial. In this case, the receiver can decode sender $2$’s message with probability 1, regardless of what sender $1$ inputs into the channel.
Case 2b: Both senders are trustworthy. Initially, the receiver can decode sender $2$’s message with probability 1. Since $D_{0}^{(2)}$ and $D_{1}^{(2)}$ do not affect part $A$ of the channel outputs when sender $2$ is trustworthy, the receiver can then decode sender $1$’s message with probability 1.
Case 2c: Sender $2$ is adversarial, while sender $1$ is trustworthy. Consider the situation where sender $1$ wants to send $m_{0}^{1}$ and transmits 00 through the channel, while sender $2$, as part of a jamming attack, transmits $2$. The channel input is now $(0,2)$, and the channel output is:
$(|0\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 0|^{A})$
When the receiver, unaware that sender $2$ is adversarial, uses $D_{0}^{(2)}$ and $D_{1}^{(2)}$ to measure sender $2$’s message, there is a $50\%$ chance that the receiver obtains $m_{0}^{2}$, with the resulting channel output after the measurement given by:
$\frac{\sqrt{D_{0}^{(2)}}(|0\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 0|^{A})\sqrt{D_{0}^{(2)}}}{\mathsf{Tr}(D_{0}^{(2)}(|0\rangle^{A}\otimes|2\rangle^{B}))}=\frac{1}{2}(|0\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 0|^{A})+\frac{1}{2}(|1\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 1|^{A})$
If instead the receiver obtains $m_{1}^{2}$, the channel output after the measurement is:
$\frac{\sqrt{D_{1}^{(2)}}(|0\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 0|^{A})\sqrt{D_{1}^{(2)}}}{\mathsf{Tr}(D_{1}^{(2)}(|0\rangle^{A}\otimes|2\rangle^{B}))}=\frac{1}{2}(|0\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 0|^{A})-\frac{1}{2}(|0\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 1|^{A})$
When the receiver subsequently uses $D_{0}^{(1)}$ and $D_{1}^{(1)}$ to measure sender $1$’s message, there is a $50\%$ chance that the receiver obtains either $m_{0}^{1}$ or $m_{1}^{1}$.
The same scenario occurs when sender $1$ transmits $m_{1}^{1}$ (i.e., when $1$ is sent through the channel) and sender $2$ jams the channel by sending $2$. In this case, the “worse order”, where the adversarial message is decoded first, results in the second code becoming ineffective due to the interference of the first POVM decoding operator. Therefore, in case $2c$, decoding the message of the adversarial sender first renders the code unreliable.
III-B Main Result
For our channel model, we should employ a separable decoding operator for each sender’s message, rather than a simultaneous decoding approach, since not all senders can be trusted. In fact, the standard methods of quantum simultaneous decoding also utilize the same technique as the Tender Operator Lemma (cf. [15], [31]). This implies that these “simultaneous decoding operators” essentially decode one message at a time, with a few exceptions (cf. [12]).
The standard coding strategy for non-symmetrizable classical-quantum arbitrarily varying channels involves constructing a random codeword (cf. [9], see also [27], [20]). This concept was further utilized in [14] to develop a code for classical-quantum arbitrarily varying multiple-access channels.
Now, assume the receiver decodes the messages in the following order: the message from sender 1 is decoded first, followed by the message from sender 2. As shown in [5], by setting $|M_{1}|<\max_{x^{n}_{2}}\chi(p_{1},B_{x^{n}_{2}})-\delta$, a random code $E_{1}=\{x^{n}_{1}(m_{1})^{\gamma_{1}},D_{m_{1}}^{\gamma_{1}}:m_{1}\in M_{1}\}$ exists such that
$\max_{x^{n}_{2}}\frac{1}{|M_{1}|}\sum_{m_{1}}\mathtt{W}((x^{n}_{1}(m_{1}),x^{n}_{2}))D_{m_{1}}^{\gamma_{1}}>1-\epsilon,$
where, for each $m_{1}$, the elements $\{x^{n}_{1}(m_{1})^{\gamma_{1}}:\gamma_{1}\}$ are permutations of a specific codeword $x^{n}_{1}(m_{1})$, and $\{D_{m_{1}}^{\gamma_{1}}:\gamma_{1}\}$ are products of permutation operators and a single decoding operator $D_{m_{1}}$.
Thus, when the receiver receives the output corresponding to the common randomness shared with sender 1, he first applies a permutation, performs the initial measurement, and then reverses the permutation before making the second measurement. Consequently, the common randomness shared with sender 1 does not affect sender 2’s coding strategy. Sender 2 only needs to consider the code $\{x^{n}_{1}(m_{1}),D_{m_{1}}:m_{1}\in M_{1}\}$, rather than every permutation of the codeword $\{x^{\gamma_{1}}_{1}(m_{1}),D_{m_{1}}^{\gamma_{1}}:m_{1}\in M_{1}\}$.
The POVM $\{D_{m_{1}}:m_{1}\in M_{1}\}$ defines a channel $C_{E_{1}}$, where
$\rho\rightarrow\sum_{m_{1}}\sqrt{D_{m_{1}}}\rho\sqrt{D_{m_{1}}}.$
Theorem III.1.
Let $\mathtt{W}$ denote a classical-quantum 2-user Byzantine multiple-access channel. Assume that the message of sender 1 is decoded first, followed by the message of sender 2. The random capacity region of $\mathtt{W}$ is given by
$R_{1}\leq\max_{p_{1}}\min_{x^{n}_{2}}I(p_{1},B_{x^{n}_{2}}),$
$R_{2}\leq\max_{p_{2}}\min_{x^{n}_{1}}I(p_{2},E_{1}B_{x^{n}_{1}}),$
where $E_{1}B_{x^{n}_{1}}$ represents the outcome of $C_{E_{1}}\circ\mathtt{W}((x^{n}_{1}(i),\cdot))$.
Proof.
The direct part of the proof follows directly from the arbitrarily varying capacity of $\mathtt{W}((x^{n}_{1}(i),\cdot))$ after the measurement of sender 1’s message (cf. [5] and [14]). It is also evident that for our coding scheme, the rate is upper-bounded by (III.1).
Now, let us consider the converse in the general case, i.e., to show that regardless of the coding strategy we choose, the order of decoding always affects the capacity. Specifically, for the message decoded second, we must always consider $C_{E_{1}}\circ\mathtt{W}$ for some $C_{E_{1}}$.
If the receiver uses a simultaneous decoding operator $\{D_{m_{1},m_{2}}^{\gamma_{1},\gamma_{2}}:m_{1},m_{2}\}$, we can always construct separable decoding operators $\{D_{m_{1}}^{\gamma_{1},\gamma_{2}}:m_{1}\}$ and $\{D_{m_{2}}^{\gamma_{1},\gamma_{2}}:m_{2}\}$ by defining
$D_{m_{1}}^{\gamma_{1},\gamma_{2}}:=\sum_{m_{2}}D_{m_{1},m_{2}}^{\gamma_{1},\gamma_{2}}\quad\text{and}\quad D_{m_{2}}^{\gamma_{1},\gamma_{2}}:=\sum_{m_{1}}D_{m_{1},m_{2}}^{\gamma_{1},\gamma_{2}}.$
Thus, without loss of generality, we may assume that the receiver uses a separable decoding strategy, decoding sender 1’s message first with a POVM $\{D_{m_{1}}^{\gamma_{1},\gamma_{2}}:m_{1}\}$. After the measurement of sender 1’s message, the channel output will result in
$\mathrm{tr}\left(\mathtt{W}((x^{n}_{1},x^{n}_{2}))D_{m_{1}}^{\gamma_{1},\gamma_{2}}\right)^{-1}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}\mathtt{W}((x^{n}_{1},x^{n}_{2}))\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}},$
with probability $\mathrm{tr}(D_{m_{1}}^{\gamma_{1},\gamma_{2}}\mathtt{W}((x^{n}_{1},x^{n}_{2})))$.
Thus, after the measurement of sender 1’s message, when the receiver begins to decode sender 2’s message, the message transmission rate cannot exceed the arbitrarily varying channel capacity of $\mathtt{W}((x^{n}_{1},\cdot))$ within the subspace defined by $D_{m_{1}}^{\gamma_{1},\gamma_{2}}$. The arbitrarily varying channel capacity of $\mathtt{W}((x^{n}_{1},\cdot))$ into this subspace is equal to
$\min_{x^{n}_{1}}\chi(p_{2},\mathrm{tr}(D_{m_{1}}^{\gamma_{1},\gamma_{2}}\mathtt{W}((x^{n}_{1},x^{n}_{2})))^{-1}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}B_{x^{n}_{1}}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}).$
Now, we define $C_{E_{1}}$ by
$C_{E_{1}}(\rho)=\sum_{m_{1}}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}\rho\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}.$
In total, the expected capacity cannot exceed
$\sum_{m_{1}}\mathrm{tr}\left(D_{m_{1}}^{\gamma_{1},\gamma_{2}}\mathtt{W}((x^{n}_{1},x^{n}_{2}))\right)\min_{x^{n}_{1}}\chi\left(p_{2},\mathrm{tr}\left(D_{m_{1}}^{\gamma_{1},\gamma_{2}}\mathtt{W}((x^{n}_{1},x^{n}_{2}))\right)^{-1}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}B_{x^{n}_{1}}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}\right),$
which simplifies to
$\min_{x^{n}_{1}}I(p_{2},E_{1}B_{x^{n}_{1}}).$
Let us now consider a Byzantine multiple-access channel with 3 senders, where at most one sender is adversarial. By Theorem III.1, we can construct reliable random codes $\{x^{\gamma_{1}}_{1}(m_{1}),D_{m_{1}}^{\gamma_{1}}:m_{1}\in M_{1}\}$, which are permutations of a code $\{x^{n}_{1}(m_{1}),D_{m_{1}}:m_{1}\in M_{1}\}$, and $\{x^{\gamma_{2}}_{2}(m_{2}),D_{m_{2}}^{\gamma_{2}}:m_{2}\in M_{2}\}$, which are permutations of a code $\{x^{n}_{2}(m_{2}),D_{m_{2}}:m_{2}\in M_{2}\}$.
We define $C_{E_{1}}$ and $C_{E_{2}}$ in the same way as in Theorem III.1, i.e.,
$C_{E_{2}}(\rho):=\sum_{m_{2}}\mathrm{tr}(D_{m_{2}}\rho)^{-1}\sqrt{D_{m_{2}}}\rho\sqrt{D_{m_{2}}}.$
Theorem III.2.
Let $\mathtt{W}$ be a classical-quantum 3-user Byzantine multiple-access channel. Assume that the receiver decodes the messages in the following order: the message of sender 1, then the message of sender 2, and finally the message of sender 3.
The capacity region of $\mathtt{W}$ is included in the following:
$R_{1}$ $\leq\min\left(\max_{p_{1}}\min_{x^{n}_{2}}I(p_{1},B_{x^{n}_{2}}),\max_{p_{1}}\min_{x^{n}_{3}}I(p_{1},B_{x^{n}_{3}})\right)$ (1)
$R_{2}$ $\leq\min\left(\max_{p_{2}}\min_{x^{n}_{1}}I(p_{2},E_{1}B_{x^{n}_{1}}),\max_{p_{2}}\min_{x^{n}_{3}}I(p_{2},B_{x^{n}_{3}})\right)$ (2)
$R_{3}$ $\leq\min\left\{\max_{p_{3}}\min_{x^{n}_{1}}I(p_{3},E_{1}B_{x^{n}_{1}}|p_{2}),\max_{p_{3}}\min_{x^{n}_{2}}I(p_{3},E_{2}B_{x^{n}_{2}}|p_{1})\right\}$ (3)
where for $i=1,2,3$, $B_{x^{n}_{i}}$ represents the channel outcome when sender $i$ sends $x^{n}_{i}$ as a jamming attack.
Proof.
The achievable rates for the messages of sender 1 and sender 2 follow from Theorem III.1. Now, consider the situation where the receiver has already performed measurements using the codes $\{x^{\gamma_{1}}_{1}(m_{1}),D_{m_{1}}^{\gamma_{1}}:m_{1}\in M_{1}\}$, which are permutations of the code $\{x^{n}_{1}(m_{1}),D_{m_{1}}:m_{1}\in M_{1}\}$, and $\{x^{\gamma_{2}}_{2}(m_{2}),D_{m_{2}}^{\gamma_{2}}:m_{2}\in M_{2}\}$, which are permutations of the code $\{x^{n}_{2}(m_{2}),D_{m_{2}}:m_{2}\in M_{2}\}$. The receiver now aims to decode the message of sender 3.
Let $\Theta:=\{x^{n}_{1}\in X^{n}_{1},x^{n}_{2}\in X^{n}_{2}\}$ be the set of uncertainty. We define an arbitrarily varying classical-quantum channel $\{\mathbf{W}(\cdot,t):t\in\Theta\}$ as follows:
$\mathbf{W}(x^{n}_{3},x^{n}_{2})$ $=\sum_{m_{1}}\sum_{m_{2}}p(x^{n}_{1}(m_{1})|x^{n}_{3})\sqrt{D_{m_{2}}}\,\mathtt{W}(x^{n}_{1}(m_{1}),x^{n}_{2},x^{n}_{3})\,\sqrt{D_{m_{2}}}$
and
$\mathbf{W}(x^{n}_{3},x^{n}_{1})$ $=\sum_{m_{2}}\sum_{m_{1}}p(x^{n}_{2}(m_{2})|x^{n}_{3})\sqrt{D_{m_{1}}}\,\mathtt{W}(x^{n}_{1},x^{n}_{2}(m_{2}),x^{n}_{3})\,\sqrt{D_{m_{1}}}\text{ .}$
Since the receiver has already measured the messages of sender 1 and sender 2, we can write
$\sum_{m_{1}}\sum_{m_{2}}\sqrt{D_{m_{1}}}\sqrt{D_{m_{2}}}\,\mathtt{W}(\cdot,\cdot,x^{n}_{3})\,\sqrt{D_{m_{1}}}\sqrt{D_{m_{2}}}$
as
$\bigoplus_{m_{1}}\left(\sum_{m_{2}}\sqrt{D_{m_{2}}}\,\sqrt{D_{m_{1}}}\,\mathtt{W}(\cdot,\cdot,x^{n}_{3})\,\sqrt{D_{m_{1}}}\,\sqrt{D_{m_{2}}}\right),$
and as
$\bigoplus_{m_{2}}\left(\sum_{m_{1}}\sqrt{D_{m_{2}}}\,\sqrt{D_{m_{1}}}\,\mathtt{W}(\cdot,\cdot,x^{n}_{3})\,\sqrt{D_{m_{1}}}\,\sqrt{D_{m_{2}}}\right).$
Thus, we have
$\chi(p_{3},\mathbf{W}(p_{3},x^{n}_{1}))$ $=S\left(\sum_{x^{n}_{3}}\sum_{m_{2}}p(x^{n}_{3})p(x^{n}_{2}(m_{2})|x^{n}_{3})\sqrt{D_{m_{2}}}\,\sqrt{D_{m_{1}}}\,\mathtt{W}(x^{n}_{1},x^{n}_{2}(m_{2}),x^{n}_{3})\,\sqrt{D_{m_{1}}}\,\sqrt{D_{m_{2}}}\right)$
$\quad-\sum_{x^{n}_{3}}p(x^{n}_{3})S\left(\sum_{m_{2}}p(x^{n}_{2}(m_{2})|x^{n}_{3})\sqrt{D_{m_{2}}}\,\sqrt{D_{m_{1}}}\,\mathtt{W}(x^{n}_{1},x^{n}_{2}(m_{2}),x^{n}_{3})\,\sqrt{D_{m_{1}}}\,\sqrt{D_{m_{2}}}\right)$
$=S\left(\sum_{x^{n}_{3}}\bigoplus_{m_{2}}p(x^{n}_{2}(m_{2}),x^{n}_{3})\sqrt{D_{m_{2}}}\,\sqrt{D_{m_{1}}}\,\mathtt{W}(x^{n}_{1},x^{n}_{2}(m_{2}),x^{n}_{3})\,\sqrt{D_{m_{1}}}\,\sqrt{D_{m_{1}}}\right)$
$\quad-\sum_{x^{n}_{3}}p(x^{n}_{3})S\left(\bigoplus_{m_{2}}p(x^{n}_{2}(m_{2})|x^{n}_{3})\sqrt{D_{m_{2}}}\,\sqrt{D_{m_{1}}}\,\mathtt{W}(x^{n}_{1},x^{n}_{2}(m_{2}),x^{n}_{3})\,\sqrt{D_{m_{1}}}\,\sqrt{D_{m_{2}}}\right)$
$=\sum_{m_{2}}p(x^{n}_{2}(m_{2}))S\left(\sum_{x^{n}_{3}}p(x^{n}_{3})p(x^{n}_{3}|x^{n}_{2}(m_{2}))\sqrt{D_{m_{1}}}\,\sqrt{D_{m_{1}}}\,\mathtt{W}(x^{n}_{1},x^{n}_{2}(m_{2}),x^{n}_{3})\,\sqrt{D_{m_{1}}}\,\sqrt{D_{m_{2}}}\right)$
$\quad-\sum_{x^{n}_{3}}\sum_{m_{2}}p(x^{n}_{3})p(x^{n}_{2}(m_{2})|x^{n}_{3})S\left(\sqrt{D_{m_{2}}}\,\sqrt{D_{m_{1}}}\,\mathtt{W}(x^{n}_{1},x^{n}_{2}(m_{2}),x^{n}_{3})\,\sqrt{D_{m_{1}}}\,\sqrt{D_{m_{2}}}\right)$
$=\sum_{m_{2}}p(x^{n}_{2}(m_{2}))S\left(\sum_{x^{n}_{3}}p(x^{n}_{3})p(x^{n}_{3}|x^{n}_{2}(m_{2}))\sqrt{D_{m_{1}}}\,\sqrt{D_{m_{1}}}\,\mathtt{W}(x^{n}_{1},x^{n}_{2}(m_{2}),x^{n}_{3})\,\sqrt{D_{m_{1}}}\,\sqrt{D_{m_{2}}}\right)$
$\quad-\sum_{m_{2}}p(x^{n}_{2}(m_{2}))\sum_{x^{n}_{3}}p(x^{n}_{3}|x^{n}_{2}(m_{2}))S\left(\sqrt{D_{m_{2}}}\,\sqrt{D_{m_{1}}}\,\mathtt{W}(x^{n}_{1},x^{n}_{2}(m_{2}),x^{n}_{3})\,\sqrt{D_{m_{1}}}\,\sqrt{D_{m_{2}}}\right)$
$=I(p_{3},E_{1}E_{2}B_{x^{n}_{1}}|p_{2}).$ (4)
Similarly, by analogous arguments, we obtain
$\chi(p_{3},\mathbf{W}(p_{3},x^{n}_{2}))=I(p_{3},E_{1}E_{2}B_{x^{n}_{2}}|p_{1}).$
Notice that when the code for the legitimate sender is reliable, then, by the Tender Operator Lemma, only the jammer’s decoding operator affects the channel output. Therefore, by [5] and [14], and if we set
$|M_{3}|$ $<\max_{p_{3}}\min_{t\in\Theta}\chi(p_{3},\mathbf{W}(p_{3},t))$
$=\min\left(\max_{p_{3}}\min_{x^{n}_{1}\in X^{n}_{1}}\chi(p_{3},\mathbf{W}(p_{3},x^{n}_{1})),\max_{p_{3}}\min_{x^{n}_{2}\in X^{n}_{2}}\chi(p_{3},\mathbf{W}(p_{3},x^{n}_{2}))\right)$
$=\min\left\{\max_{p_{3}}\min_{x^{n}_{1}\in X^{n}_{1}}I(p_{3},E_{1}B_{x^{n}_{1}}|p_{2}),\max_{p_{3}}\min_{x^{n}_{2}\in X^{n}_{2}}I(p_{3},E_{2}B_{x^{n}_{2}}|p_{1})\right\},$
for sufficiently large $n$, there exists a reliable random code
$\left\{\left\{x^{\gamma_{3}}_{3}(m_{3}),D_{m_{3}}^{\gamma_{3}}:m_{3}\in M_{3}\right\}:\gamma_{3}\right\}$
a permutation of a code $\{x^{n}_{3}(m_{3}),D_{m_{3}}:m_{3}\in M_{3}\}$, such that for any positive $\epsilon$ and sufficiently large $n$,
$1-\epsilon$ $<\max_{x^{n}_{1}\in X^{n}_{1}}\frac{1}{|\Lambda_{3}|}\sum_{\gamma_{3}\in\Lambda_{3}}\mathrm{tr}\left(\sum_{m_{3}}p(m_{3})\mathbf{W}(x^{\gamma_{3}}_{3}(m_{3}),x^{n}_{1})D_{m_{3}}^{\gamma_{3}}\right)$
$=\max_{x^{n}_{1}\in X^{n}_{1}}\frac{1}{|\Lambda_{3}|}\sum_{\gamma_{3}\in\Lambda_{3}}\mathrm{tr}\left(\sum_{m_{3}}\sum_{m_{2}}p(m_{3})p(x^{n}_{2}(m_{2})|x^{\gamma_{3}}_{3}(m_{3}))C_{E_{1}}\mathtt{W}(x^{n}_{1},x^{n}_{2}(m_{2}),x^{\gamma_{3}}_{3}(m_{3}))D_{m_{3}}^{\gamma_{3}}\right),$
and
$1-\epsilon$ $<\max_{x^{n}_{2}\in X^{n}_{2}}\frac{1}{|\Lambda_{3}|}\sum_{\gamma_{3}\in\Lambda_{3}}\mathrm{tr}\left(\sum_{m_{3}}p(m_{3})\mathbf{W}(x^{\gamma_{3}}_{3}(m_{3}),x^{n}_{2})D_{m_{3}}^{\gamma_{3}}\right)$
$=\max_{x^{n}_{2}\in X^{n}_{2}}\frac{1}{|\Lambda_{3}|}\sum_{\gamma_{3}\in\Lambda_{3}}\mathrm{tr}\left(\sum_{m_{3}}\sum_{m_{1}}p(m_{3})p(x^{n}_{1}(m_{1})|x^{\gamma_{3}}_{3}(m_{3}))C_{E_{2}}\mathtt{W}(x^{n}_{1}(m_{1}),x^{n}_{2},x^{\gamma_{3}}_{3}(m_{3}))D_{m_{3}}^{\gamma_{3}}\right).$
Let us first consider the case where sender 1 is the adversarial one and sends $x^{n}_{1}$ as a jamming attack. The legitimate senders, sender 2 and sender 3, send $x^{\gamma_{2}}_{2}$ and $x^{\gamma_{3}}_{3}$, respectively. The receiver initially performs a measurement with $\{D_{m_{1}}^{\gamma_{1}}:m_{1}\}$, and then reverses the permutation $\gamma_{1}$. The resulting channel output occurs with probability
$\mathrm{tr}\left(D_{m_{1}}\mathtt{W}\left((x^{n}_{1},x^{\gamma_{2}}_{2}(m_{2}),x^{\gamma_{3}}_{3}(m_{3}))\right)\right),$
and the output is given by
$\mathrm{tr}\left(D_{m_{1}}\mathtt{W}\left((x^{n}_{1},x^{\gamma_{2}}_{2}(m_{2}),x^{\gamma_{3}}_{3}(m_{3}))\right)\right)^{-1}\sqrt{D_{m_{1}}}\mathtt{W}\left((x^{n}_{1},x^{\gamma_{2}}_{2}(m_{2}),x^{\gamma_{3}}_{3}(m_{3}))\right)\sqrt{D_{m_{1}}}.$
Next, the receiver performs a measurement with $\{D_{m_{2}}^{\gamma_{2}}:m_{2}\}$ and then reverses the permutation $\gamma_{2}$. By the Tender Operator Lemma, the channel output changes negligibly. Afterward, he performs a measurement with $\{D_{m_{3}}^{\gamma_{3}}:m_{3}\}$ and obtains $M_{3}$ with probability
$\frac{1}{|\Lambda_{3}|}\sum_{\gamma_{3}}\mathrm{tr}\left(D_{m_{1}}\mathtt{W}\left((x^{n}_{1},x^{\gamma_{2}}_{2}(m_{2}),x^{\gamma_{3}}_{3}(m_{3}))\right)\right)^{-1}\mathrm{tr}\left(C_{E_{1}}\circ\mathtt{W}\left((x^{n}_{1},x^{\gamma_{2}}_{2}(m_{2}),x^{\gamma_{3}}_{3}(m_{3}))\right)D_{m_{3}}^{\gamma_{3}}\right)^{-1}.$
Thus, the expected value of correctly decoding $m_{3}$ is
$\frac{1}{|\Lambda_{3}|}\sum_{\gamma_{3}\in\Lambda_{3}}\mathrm{tr}\left(\sum_{m_{3}}\sum_{m_{1}}p(m_{3})p(x^{n}_{1}(m_{1})|x^{\gamma_{3}}_{3}(m_{3}))C_{E_{2}}\mathtt{W}\left((x^{n}_{1}(m_{1}),x^{n}_{2},x^{\gamma_{3}}_{3}(m_{3}))\right)D_{m_{3}}^{\gamma_{3}}\right).$
Therefore, this code is a reliable random code for sender 3 to transmit his message through $\mathtt{W}$.
The same technique can be extended to a Byzantine multiple-access channel with $k$ senders and at most one adversarial sender:
Corollary III.3.
Let $\mathtt{W}$ be a classical-quantum $k$-user Byzantine multiple-access channel with at most one adversarial sender. Assume that the receiver decodes the messages in the order: sender 1, sender 2, …, sender $k$. For each sender $i\in\{1,\dots,k\}$, we denote the arbitrarily varying classical-quantum channel that maps the message of sender $i$ into the output state after having already measured the messages of senders $11,\dots,i-$ 1 by $\mathtt{W}_{i}^{j}$, where $x^{n}_{j}$ and $\gamma_{j}$ represent the jamming attack by the adversarial sender $j\in\{1,\dots,i-1\}$.
The random capacity region of $\mathtt{W}$ is included in
$R_{i}\leq\min\Bigl{(}\max_{p_{i}}\min_{x^{n}_{j}:j<i}I(p_{i},E_{j}B_{x^{n}_{j}}|\{p_{h}:h\in\{1,\dots,i-1\}\setminus j\}),$ (5)
$\max_{p_{i}}\min_{x^{n}_{j}:j>i}I(p_{i},B_{x^{n}_{j}}|\{p_{h}:h\in\{1,\dots,i-1\}\})\Bigr{)}\text{ .}$ (6)
Proof.
Similar to the proof of Theorem III.2, we define
$\Theta_{i}:=\{x^{n}_{j}:j\in\{1,\dots,k\}\setminus i\}$
and an arbitrarily varying classical-quantum channel
$\{\mathbf{W}(\cdot,t):t\in\Theta_{i}\}:X_{i}\rightarrow X_{1}\times\cdots\times X_{i-1}\otimes\mathcal{S}(A)$
defined by
$\mathbf{W}(x^{n}_{i},x^{n}_{j})$
$=\sum_{m_{1},\dots,m_{k}}p(x^{n}_{1}(m_{1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{i-1}(m_{i-1}),x^{n}_{i+1}(m_{i+1}),\dots,x^{n}_{k}(m_{k})|x^{n}_{i})$
$\quad\cdot x^{n}_{1}(m_{1})\otimes\cdots\otimes x^{n}_{i-1}(m_{i-1})$
$\quad\otimes\sqrt{D_{m_{j}}}\mathtt{W}((x^{n}_{1}(m_{1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j},x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{i-1}(m_{i-1}),x^{n}_{i},x^{n}_{i+1}(m_{i+1}),\dots,x^{n}_{k}(m_{k})))\sqrt{D_{m_{j}}}$
for $j<i$, and
$\mathbf{W}(x^{n}_{i},x^{n}_{j})$ $=\sum_{m_{1},\dots,m_{k}}p(x^{n}_{1}(m_{1}),\dots,x^{n}_{i-1}(m_{i-1}),x^{n}_{i+1}(m_{i+1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k})|x^{n}_{i})$
$\quad\cdot x^{n}_{1}(m_{1})\otimes\cdots\otimes x^{n}_{i-1}(m_{i-1})$
$\quad\otimes\mathtt{W}((x^{n}_{1}(m_{1}),\dots,x^{n}_{i-1}(m_{i-1}),x^{n}_{i},x^{n}_{i+1}(m_{i+1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j},x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k})))$
for $j>i$.
When
$|M_{3}|<\min_{x^{n}_{j}:j<i}I(p_{i},E_{j}B_{x^{n}_{j}}|\{p_{h}:h\in\{1,\dots,i-1\}\setminus j\})$
and
$|M_{3}|<\min_{x^{n}_{j}:j>i}I(p_{i},B_{x^{n}_{j}}|\{p_{h}:h\in\{1,\dots,i-1\}\}),$
then by [5] and [14], there exists a reliable random code for $\{\mathbf{W}(\cdot,t):t\in\Theta_{i}\}$.
The receiver now applies a similar strategy to the quantum multiple access channel coding in [34]. We assume that, after having already applied the measurements for the messages of senders 1 through $i-1$, the channel output results in
$\mathrm{tr}\Bigl{(}\mathtt{W}((x^{n}_{1}(m_{1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j},x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k}))\prod_{h=1}^{i-1}D_{m_{h}}\Bigr{)}^{-1}$
$\sqrt{D_{m_{1}}}\cdot\cdots\cdot\sqrt{D_{m_{i-1}}}\mathtt{W}((x^{n}_{1}(m_{1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j},x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k})))\sqrt{D_{m_{i-1}}}\cdot\cdots\cdot\sqrt{D_{m_{1}}}$
for some $m_{1},\dots,m_{i-1}$ with probability
$\mathrm{tr}\Bigl{(}\mathtt{W}((x^{n}_{1}(m_{1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j},x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k}))\prod_{h=1}^{i-1}D_{m_{h}}\Bigr{)}.$
Now the receiver uses
$\left\{\mathrm{tr}\left(\sum_{m_{i}}D_{m_{i}}^{(m_{1},\dots,m_{i-1})}\right)^{-1}D_{m_{i}}^{(m_{1},\dots,m_{i-1})}:m_{i}\right\}$
to decode the message of sender $i$. Let us first consider the adversarial sender $j$ when $j>i$. By the Tender Operator Lemma, the channel output
$\mathrm{tr}\left(\mathtt{W}((x^{n}_{1}(m_{1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j},x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k}))\prod_{h=1}^{i-1}D_{m_{h}}\right)^{-1}$
$\sqrt{D_{m_{1}}}\cdot\dots\cdot\sqrt{D_{m_{i-1}}}\mathtt{W}((x^{n}_{1}(m_{1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j},x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k})))$
$\sqrt{D_{m_{i-1}}}\cdot\dots\cdot\sqrt{D_{m_{1}}}$
after decoding the messages of senders 1 to $i-1$ is close to the original channel output
$\mathtt{W}((x^{n}_{1}(m_{1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j},x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k}))).$
The expected value of a successful decoding is thus:
$\min_{x^{n}_{j}}\frac{1}{|\{m_{i}\}|}\sum_{m_{1},\dots,m_{k}}p(x^{n}_{1}(m_{1}),\dots,x^{n}_{k}(m_{k}))$
×$\quad\times\mathrm{tr}\left(\mathtt{W}((x^{n}_{1}(m_{1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j},x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k}))D_{m_{i}}^{(m_{1},\dots,m_{i-1})}\right)$
$=\min_{x^{n}_{j}}\frac{1}{|\{m_{i}\}|}\sum_{m_{i}}p(x^{n}_{i}(m_{i}))$
×$\quad\times\sum_{m_{1},\dots,m_{k}}p(x^{n}_{1}(m_{1}),\dots,x^{n}_{i-1}(m_{i-1}),x^{n}_{i+1}(m_{i+1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k})|x^{n}_{i}(m_{i}))$
×$\quad\times\mathrm{tr}(\Bigl{(}\sqrt{D_{m_{1}}}\cdot\dots\cdot\sqrt{D_{m_{i-1}}}\mathtt{W}((x^{n}_{1}(m_{1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j},x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k})))$
×$\quad\left.\times\sqrt{D_{m_{i-1}}}\cdot\dots\cdot\sqrt{D_{m_{1}}}D_{m_{i}}^{(m_{1},\dots,m_{i-1})}\Bigr{)}\right)-\epsilon$
$=\min_{x^{n}_{j}}\frac{1}{|\{m_{i}\}|}\mathrm{tr}\left(\mathbf{W}(x^{n}_{i}(m_{i}),x^{n}_{j})D_{m_{i}}\right)-\epsilon$
for a positive $\epsilon$ that can be made arbitrarily close to zero for sufficiently large $n$. Similarly, the expected value of a successful decoding for the adversarial sender $j$ when $j<i$, for sufficiently large $n$, is arbitrarily close to:
$\quad\min_{x^{n}_{j}}\frac{1}{|\{m_{i}\}|}\sum_{m_{1},\dots,m_{k}}p(x^{n}_{1}(m_{1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k}))$
×$\quad\times\mathrm{tr}\Bigl{(}\sqrt{D_{m_{j}}}\mathtt{W}((x^{n}_{1}(m_{1}),\dots,x^{n}_{i-1}(m_{i-1}),x^{n}_{i}(m_{i}),x^{n}_{i+1}(m_{i+1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j},x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k})))$
$\quad\sqrt{D_{m_{j}}}D_{m_{i}}^{(m_{1},\dots,m_{i-1})}\Bigr{)}$
$=\min_{x^{n}_{j}}\frac{1}{|\{m_{i}\}|}\sum_{m_{i}}p(x^{n}_{i}(m_{i}))$
×$\quad\times\sum_{m_{1},\dots,m_{k}}p(x^{n}_{1}(m_{1}),\dots,x^{n}_{i-1}(m_{i-1}),x^{n}_{i+1}(m_{i+1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k})|x^{n}_{i}(m_{i}))$
×$\quad\times\mathrm{tr}(\Bigl{(}x^{n}_{1}(m_{1})\otimes\dots\otimes x^{n}_{i-1}(m_{i-1})$
$\quad\sqrt{D_{m_{j}}}\mathtt{W}((x^{n}_{1}(m_{1}),\dots,x^{n}_{i-1}(m_{i-1}),x^{n}_{i}(m_{i}),x^{n}_{i+1}(m_{i+1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j},x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k})))$
×$\quad\left.\times\sqrt{D_{m_{j}}}\Bigr{)}D_{m_{i}}\right)$
$=\min_{x^{n}_{j}}\frac{1}{|\{m_{i}\}|}\mathrm{tr}\left(\mathbf{W}(x^{n}_{i}(m_{i}),x^{n}_{j})D_{m_{i}}\right)$
This shows that the receiver can utilize the part of the code corresponding to $m_{1},\dots,m_{i-1}$ for the channel $\{\mathbf{W}(\cdot,t):t\in\Theta_{i}\}$ as an arbitrarily varying channel code in order to decode the message of sender $i$.
As stated in Section III-A, the channel output can be altered by the decoding operator. In this section, we provide an example of a family of channels that are immune to adversarial decoding operators. We introduce a new symmetrizable condition as follows:
Definition III.4.
We say that the arbitrarily varying classical-quantum channel $\{\mathbf{W}(\cdot,t):t\in\theta\}$ is orthogonally symmetrizable if there exists a parametrized set of distributions $\{\tau(\cdot\mid x):x\in X\}$ on $\theta$ such that for all $x,x^{\prime}\in X$,
$\mathrm{tr}\left(\left(\sum_{t\in\theta}\tau(t\mid x)\mathbf{W}(x^{\prime},t)\right)\left(\sum_{t\in\theta}\tau(t\mid x^{\prime})\mathbf{W}(x,t)\right)\right)>0.$
It is clear that non-orthogonal symmetrizability is a very strong assumption, and very few channels will satisfy this condition. However, channels that are not orthogonally symmetrizable are still not affected by hostile decoding operators. We demonstrate this result for a $2$-user Byzantine multiple access channel, though this result can also be extended to a $k$-user Byzantine multiple access channel.
Corollary III.5.
Let $\mathtt{W}$ be a classical-quantum $2$-user Byzantine multiple access channel. Assume that the message of sender $1$ is to be decoded first, followed by the message of sender $2$. If $x^{n}_{1}\rightarrow\mathtt{W}((x^{n}_{1},x^{n}_{2}))$ is not symmetrizable, and $x^{n}_{2}\rightarrow\mathtt{W}((x^{n}_{1},x^{n}_{2}))$ is not orthogonally symmetrizable, then the capacity region of $\mathtt{W}$ is given by
$R_{1}$ $\leq\max_{p_{1}}\min_{x^{n}_{2}}I(p_{1},B_{x^{n}_{2}})$ (7)
$R_{2}$ $\leq\max_{p_{2}}\min_{x^{n}_{1}}I(p_{2},B_{x^{n}_{1}})\text{ .}$ (8)
Proof.
The receiver first performs a measurement using an arbitrarily varying decoding operator set $\{D^{\gamma_{1}}_{m_{1}}:m_{1}\in M_{1}\}$, which is constructed from the typical space projectors of the presumed code words of sender 1 for the arbitrarily varying channel $x^{n}_{1}\rightarrow\mathtt{W}((x^{n}_{1},x^{n}_{2}))$. If the size is smaller than $\min_{x^{n}_{2}}I(p_{1},B_{x^{n}_{2}})$, the receiver can almost certainly decode the message of sender 1, even when sender 2 acts as a jammer.
Next, we consider the case where sender 1 is adversarial, sending a sequence $x_{1}^{n}$ as an attack, while sender 2 transmits $x_{2}(m_{2})^{\gamma_{2}}$ for $m_{2}\in M_{2}$. After the first measurement, the receiver then performs a second measurement using arbitrarily varying decoding operators $\{D^{\gamma_{2}}_{(m_{1},m_{2})}:m_{1}\in M_{1},m_{2}\in M_{2}\}$ for the arbitrarily varying channel $(x^{n}_{2})\rightarrow\mathtt{W}((x^{n}_{1},x^{n}_{2}))$ with the knowledge of $m_{1}$ (cf. [34]).
When $x^{n}_{2}\rightarrow\mathtt{W}((x^{n}_{1},x^{n}_{2}))$ is not orthogonally symmetrizable, we have that $\mathtt{W}((x^{n}_{1},x^{n}_{2}(m_{2})^{\gamma_{2}}))\perp\mathtt{W}((x^{n}_{1}(m_{1}),x^{n}_{2}(m_{2}^{\prime})^{\gamma_{2}}))$ for all $m_{2}^{\prime}\neq m_{2}$. Thus, we can construct a set $\{D^{\gamma_{2}}_{(m_{1},m_{2})}:m_{2}\in M_{2}\}$ such that
$\mathrm{tr}\left(\mathtt{W}((x^{n}_{1},x^{n}_{2}(m_{2})))D_{m_{1}}\right)^{-1}\sqrt{D_{m_{1}}}\mathtt{W}((x^{n}_{1},x^{n}_{2}(m_{2})))\sqrt{D_{m_{1}}}D^{\gamma_{2}}_{(m_{1},m_{2}^{\prime})}<\epsilon$
for all $m_{2}^{\prime}\neq m_{2}$. This ensures that the error probability of the code remains small after the first measurement on the channel output.
IV Conclusion and Future Directions
In this work, we explored the impact of adversarial senders on Byzantine multiple-access classical-quantum channels. We demonstrated that a $2$-user Byzantine multiple-access channel, which may be a straightforward model in the classical setting, can exhibit non-trivial behavior in the quantum context. Specifically, we derived the capacity region for the $2$-user Byzantine multiple-access classical-quantum channels. Additionally, we established an achievable rate region for the $k$-user Byzantine multiple-access classical-quantum channels when $k\geq 3$.
A general upper bound for the rate region with more than three senders remains an open problem. Furthermore, an important challenge lies in applying advanced classical techniques, such as adversarial sender identification from [29] and [30], to quantum channels. The main difficulty, as mentioned earlier, is that for quantum arbitrarily varying channels, only random coding techniques are currently known.
Acknowledgments
The authors acknowledge the support of the German Federal Ministry of Education and Research (BMBF) under several national initiatives: the 6G Communication Systems initiative through the research hub 6G-life (grant 16KISK263), the QuaPhySI (Quantum Physical Layer Service Integration) initiative (grant 16KIS2234), the “QTOK - Quantum Tokens for Secure Authentication in Theory and Practice” initiative (grant 16KISQ038), the QR.N initiative (grant 16KIS2196), the QTREX initiative (grant 16KISR038), the QUIET initiative (grant 16KISQ0170), and the QD-CAMNetz initiative (grant 16ISQ169).

### v2

@@
 In this work, we investigate a compelling scenario: a multiple-access channel with two transmitters and one receiver, where one transmitter deviates from the protocol and acts dishonestly.
-To address this challenge, we introduce the Byzantine multiple-access classical-quantum channel and derive an achievable communication rate for this adversarial setting.
+To address this challenge, we introduce the Byzantine multiple-access classical-quantum channel and derive the communication region of rate for this adversarial setting.
 I Introduction
@@
 As a result, many of the sophisticated techniques from [29] cannot be directly extended to quantum Byzantine multiple-access channels.
-Similarly, the identification technique proposed in [30] is not applicable: while a sequence of classical symbols can be easily verified against a set of expected legal channel outputs, a “false” quantum state may still have a nonzero probability of being accepted by the decoding operators for legitimate messages.
+Similarly, the identification technique proposed in [30] is not applicable: while a sequence of classical symbols can be easily verified against a set of expected legal channel outputs, a "false" quantum state may still have a nonzero probability of being accepted by the decoding operators for legitimate messages.
 To overcome these challenges, we must develop a novel coding approach.
@@
 $\chi(P;\mathtt{V}):=S\left(\sum_{x\in X}P(x)\mathtt{V}(x)\right)-S(\mathtt{V}|P).$
-Notice that when we choose an “optimal” set $\Phi:=\{\rho_{x}:x\in X\}$ (in the sense that it maximizes the quantum mutual information and the Holevo quantity), we have (cf. [21])
+Notice that when we choose an "optimal" set $\Phi:=\{\rho_{x}:x\in X\}$ (in the sense that it maximizes the quantum mutual information and the Holevo quantity), we have (cf. [21])
 $I(P;\mathtt{V})=\chi(P;\mathtt{V})\quad\text{.}$
@@
 $\sum_{t\in\Theta}\tau(t\mid x)\mathbf{W}(x^{\prime},t)=\sum_{t\in\Theta}\tau(t\mid x^{\prime})\mathbf{W}(x,t).$
+For a probability distribution $p$ on $\Theta$, we denote $\mathbf{W}(\cdot,p):=\sum_{t}\mathbf{W}(\cdot,t)$.
 Definition II.2.
@@
 III Message Transmission over Quantum Byzantine Multiple Access Channel
-III-A Necessity of Decoding Order
+III-A Impact of Decoding Order
 In the classical multiple-access channel scenario, it is evident that the order in which the messages of the senders are decoded does not matter.
@@
 Therefore, in this scenario, it is possible that if the first sender is adversarial, the first decoding operator may alter the channel outputs in such a way that the second message becomes undecodable.
-The following simple example illustrates this point:
+The following simple one-shot example illustrates this point:
 We define the input alphabets as $X_{1}:=\{0,1\}$ and $X_{2}:=\{0,1,2\}$. Let $A\otimes B$ be a six-dimensional Hilbert space spanned by the orthonormal basis:
@@
 $(|0\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 0|^{A})$
-When the receiver, unaware that sender $2$ is adversarial, uses $D_{0}^{(2)}$ and $D_{1}^{(2)}$ to measure sender $2$’s message, there is a $50\%$ chance that the receiver obtains $m_{0}^{2}$, with the resulting channel output after the measurement given by:
+When the receiver, unaware that sender $2$ is adversarial, uses $D_{0}^{(2)}$ and $D_{1}^{(2)}$ to measure sender $2$’s message, there is a $50\%$ chance that the receiver obtains either $m_{0}^{2}$ or $m_{1}^{2}$.
+When he obtains $m_{0}^{2}$, then the resulting channel output after the measurement is given by:
 $\frac{\sqrt{D_{0}^{(2)}}(|0\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 0|^{A})\sqrt{D_{0}^{(2)}}}{\mathsf{Tr}(D_{0}^{(2)}(|0\rangle^{A}\otimes|2\rangle^{B}))}=\frac{1}{2}(|0\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 0|^{A})+\frac{1}{2}(|1\rangle^{A}\otimes|2\rangle^{B})(\langle 2|^{B}\otimes\langle 1|^{A})$
@@
 The same scenario occurs when sender $1$ transmits $m_{1}^{1}$ (i.e., when $1$ is sent through the channel) and sender $2$ jams the channel by sending $2$.
-In this case, the “worse order”, where the adversarial message is decoded first, results in the second code becoming ineffective due to the interference of the first POVM decoding operator.
+In this case, the "worse order", where the adversarial message is decoded first, results in the second code becoming ineffective due to the interference of the first POVM decoding operator.
 Therefore, in case $2c$, decoding the message of the adversarial sender first renders the code unreliable.
@@
 The random capacity region of $\mathtt{W}$ is given by
-$R_{1}\leq\max_{p_{1}}\min_{x^{n}_{2}}I(p_{1},B_{x^{n}_{2}}),$
-$R_{2}\leq\max_{p_{2}}\min_{x^{n}_{1}}I(p_{2},E_{1}B_{x^{n}_{1}}),$
-where $E_{1}B_{x^{n}_{1}}$ represents the outcome of $C_{E_{1}}\circ\mathtt{W}((x^{n}_{1}(i),\cdot))$.
+$R_{1}\leq\max_{p_{1}}\min_{p_{2}}I(p_{1},B),$
+$R_{2}\leq\max_{p_{2}}\min_{p_{1}}I(p_{2},E_{1}B),$
+where $B$ represents the outcome of $\mathtt{W}((p_{1},p_{2})$. and $E_{1}B$ represents the outcome of $C_{E_{1}}\circ\mathtt{W}(p_{1},p_{2})$.
 Proof.
@@
 with probability $\mathrm{tr}(D_{m_{1}}^{\gamma_{1},\gamma_{2}}\mathtt{W}((x^{n}_{1},x^{n}_{2})))$.
-Thus, after the measurement of sender 1’s message, when the receiver begins to decode sender 2’s message, the message transmission rate cannot exceed the arbitrarily varying channel capacity of $\mathtt{W}((x^{n}_{1},\cdot))$ within the subspace defined by $D_{m_{1}}^{\gamma_{1},\gamma_{2}}$.
-The arbitrarily varying channel capacity of $\mathtt{W}((x^{n}_{1},\cdot))$ into this subspace is equal to
-$\min_{x^{n}_{1}}\chi(p_{2},\mathrm{tr}(D_{m_{1}}^{\gamma_{1},\gamma_{2}}\mathtt{W}((x^{n}_{1},x^{n}_{2})))^{-1}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}B_{x^{n}_{1}}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}).$
+Thus, after the measurement of sender 1’s message, when the receiver begins to decode sender 2’s message, the message transmission rate cannot exceed the arbitrarily varying channel capacity of $\mathtt{W}(p_{1},p_{2})$ within the subspace defined by $D_{m_{1}}^{\gamma_{1},\gamma_{2}}$.
+The arbitrarily varying channel capacity of $\mathtt{W}((p_{1},\cdot))$ into this subspace is equal to
+$\min_{x^{n}_{1}}\chi(p_{2},\mathrm{tr}(D_{m_{1}}^{\gamma_{1},\gamma_{2}}\mathtt{W}((p_{1},x^{n}_{2})))^{-1}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}B_{p_{1},p_{2}}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}).$
 Now, we define $C_{E_{1}}$ by
@@
 In total, the expected capacity cannot exceed
-$\sum_{m_{1}}\mathrm{tr}\left(D_{m_{1}}^{\gamma_{1},\gamma_{2}}\mathtt{W}((x^{n}_{1},x^{n}_{2}))\right)\min_{x^{n}_{1}}\chi\left(p_{2},\mathrm{tr}\left(D_{m_{1}}^{\gamma_{1},\gamma_{2}}\mathtt{W}((x^{n}_{1},x^{n}_{2}))\right)^{-1}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}B_{x^{n}_{1}}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}\right),$
+$\sum_{m_{1}}\mathrm{tr}\left(D_{m_{1}}^{\gamma_{1},\gamma_{2}}\mathtt{W}((x^{n}_{1},x^{n}_{2}))\right)\min_{x^{n}_{1}}\chi\left(p_{2},\mathrm{tr}\left(D_{m_{1}}^{\gamma_{1},\gamma_{2}}\mathtt{W}((x^{n}_{1},x^{n}_{2}))\right)^{-1}\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}B\sqrt{D_{m_{1}}^{\gamma_{1},\gamma_{2}}}\right),$
 which simplifies to
-$\min_{x^{n}_{1}}I(p_{2},E_{1}B_{x^{n}_{1}}).$
+$\min_{p_{1}}I(p_{2},E_{1}B).$
 Let us now consider a Byzantine multiple-access channel with 3 senders, where at most one sender is adversarial.
@@
 Assume that the receiver decodes the messages in the following order: the message of sender 1, then the message of sender 2, and finally the message of sender 3.
-The capacity region of $\mathtt{W}$ is included in the following:
-$R_{1}$ $\leq\min\left(\max_{p_{1}}\min_{x^{n}_{2}}I(p_{1},B_{x^{n}_{2}}),\max_{p_{1}}\min_{x^{n}_{3}}I(p_{1},B_{x^{n}_{3}})\right)$ (1)
-$R_{2}$ $\leq\min\left(\max_{p_{2}}\min_{x^{n}_{1}}I(p_{2},E_{1}B_{x^{n}_{1}}),\max_{p_{2}}\min_{x^{n}_{3}}I(p_{2},B_{x^{n}_{3}})\right)$ (2)
-$R_{3}$ $\leq\min\left\{\max_{p_{3}}\min_{x^{n}_{1}}I(p_{3},E_{1}B_{x^{n}_{1}}|p_{2}),\max_{p_{3}}\min_{x^{n}_{2}}I(p_{3},E_{2}B_{x^{n}_{2}}|p_{1})\right\}$ (3)
-where for $i=1,2,3$, $B_{x^{n}_{i}}$ represents the channel outcome when sender $i$ sends $x^{n}_{i}$ as a jamming attack.
+The capacity region of $\mathtt{W}$ is the following:
+$R_{1}$ $\leq\min\left(\max_{p_{1}}\min_{p_{2}}I(p_{1},B),\max_{p_{1}}\min_{p_{3}}I(p_{1},B)\right)$ (1)
+$R_{2}$ $\leq\min\left(\max_{p_{2}}\min_{p_{1}}I(p_{2},E_{1}B),\max_{p_{2}}\min_{p_{3}}I(p_{2},B)\right)$ (2)
+$R_{3}$ $\leq\min\left\{\max_{p_{3}}\min_{p_{1}}I(p_{3},E_{1}B|p_{2}),\max_{p_{3}}\min_{p_{2}}I(p_{3},E_{2}B|p_{1})\right\}$ (3)
 Proof.
-The achievable rates for the messages of sender 1 and sender 2 follow from Theorem III.1.
+The capacity region for the messages of sender 1 and sender 2 follow from Theorem III.1.
 Now, consider the situation where the receiver has already performed measurements using the codes $\{x^{\gamma_{1}}_{1}(m_{1}),D_{m_{1}}^{\gamma_{1}}:m_{1}\in M_{1}\}$, which are permutations of the code $\{x^{n}_{1}(m_{1}),D_{m_{1}}:m_{1}\in M_{1}\}$, and $\{x^{\gamma_{2}}_{2}(m_{2}),D_{m_{2}}^{\gamma_{2}}:m_{2}\in M_{2}\}$, which are permutations of the code $\{x^{n}_{2}(m_{2}),D_{m_{2}}:m_{2}\in M_{2}\}$.
@@
 $\quad-\sum_{m_{2}}p(x^{n}_{2}(m_{2}))\sum_{x^{n}_{3}}p(x^{n}_{3}|x^{n}_{2}(m_{2}))S\left(\sqrt{D_{m_{2}}}\,\sqrt{D_{m_{1}}}\,\mathtt{W}(x^{n}_{1},x^{n}_{2}(m_{2}),x^{n}_{3})\,\sqrt{D_{m_{1}}}\,\sqrt{D_{m_{2}}}\right)$
-$=I(p_{3},E_{1}E_{2}B_{x^{n}_{1}}|p_{2}).$ (4)
+$=I(p_{3},E_{1}E_{2}B|p_{2}).$ (4)
 Similarly, by analogous arguments, we obtain
-$\chi(p_{3},\mathbf{W}(p_{3},x^{n}_{2}))=I(p_{3},E_{1}E_{2}B_{x^{n}_{2}}|p_{1}).$
+$\chi(p_{3},\mathbf{W}(p_{3},x^{n}_{2}))=I(p_{3},E_{1}E_{2}B|p_{1}).$
+Denote the set of probability distributions on $\Theta$ by $P(\Theta)$.
 Notice that when the code for the legitimate sender is reliable, then, by the Tender Operator Lemma, only the jammer’s decoding operator affects the channel output.
 Therefore, by [5] and [14], and if we set
-$|M_{3}|$ $<\max_{p_{3}}\min_{t\in\Theta}\chi(p_{3},\mathbf{W}(p_{3},t))$
-$=\min\left(\max_{p_{3}}\min_{x^{n}_{1}\in X^{n}_{1}}\chi(p_{3},\mathbf{W}(p_{3},x^{n}_{1})),\max_{p_{3}}\min_{x^{n}_{2}\in X^{n}_{2}}\chi(p_{3},\mathbf{W}(p_{3},x^{n}_{2}))\right)$
-$=\min\left\{\max_{p_{3}}\min_{x^{n}_{1}\in X^{n}_{1}}I(p_{3},E_{1}B_{x^{n}_{1}}|p_{2}),\max_{p_{3}}\min_{x^{n}_{2}\in X^{n}_{2}}I(p_{3},E_{2}B_{x^{n}_{2}}|p_{1})\right\},$
+$|M_{3}|$ $<\max_{p_{3}}\min_{p_{t}\in P(\Theta)}\chi(p_{3},\mathbf{W}(p_{3},p_{t}))$
+$=\min\left(\max_{p_{3}}\min_{p_{1}}\chi(p_{3},\mathbf{W}(p_{3},p_{1})),\max_{p_{3}}\min_{p_{2}}\chi(p_{3},\mathbf{W}(p_{3},p_{2}))\right)$
+$=\min\left\{\max_{p_{3}}\min_{p_{1}}I(p_{3},E_{1}B|p_{2}),\max_{p_{3}}\min_{p_{2}}I(p_{3},E_{2}B|p_{1})\right\},$
 for sufficiently large $n$, there exists a reliable random code
@@
 Therefore, this code is a reliable random code for sender 3 to transmit his message through $\mathtt{W}$.
+For the converse for the message of sender 3, we may, without loss of generality, assume that the receiver uses a separable decoding strategy, as stated in the proof of Thoerem III.1 .
+We consider now the case when sender 3 sends $x^{\gamma_{3}}_{3}(m_{3})$ through the channel.
+When sender 1 is adversarial and sends $x^{n}_{1}$ as jamming attack, bthen by the Tender Operator Lemma, the channel output will be sufficiently closed to
+$\sum_{m_{2}}p(x^{n}_{1}(m_{1})|x^{\gamma_{3}}_{3}(m_{3}))\sqrt{D_{m_{1}}}\,\mathtt{W}(x^{n}_{1}(m_{1}),x^{n}_{2},x^{\gamma_{3}}_{3}(m_{3}))\,\sqrt{D_{m_{1}}}$
+with probability $\mathrm{tr}\biggl{(}\sum_{m_{2}}p(x^{n}_{1}(m_{1})|x^{\gamma_{3}}_{3}(m_{3}))$ $\sqrt{D_{m_{1}}}$ $\mathtt{W}(x^{n}_{1}(m_{1}),x^{n}_{2},x^{\gamma_{3}}_{3}(m_{3}))$ $\sqrt{D_{m_{1}}}\biggr{)}$.
+Thus, the expectetd value of the channel output will be closed to the arbitrarily varying classical-quantum channel output $\mathbf{W}(x^{\gamma_{3}}_{3}(m_{3}),x^{n}_{1}))$, When sender 2 is adversarial and sends $x^{n}_{2}$ as jamming attack, the expectetd value of the channel output will be closed to the arbitrarily varying classical-quantum channel output $\mathbf{W}(x^{\gamma_{3}}_{3}(m_{3}),x^{n}_{2}))$.
+By [5] and [14], the rate for the message of sender 3 cannot exceed
+$\max_{p_{3}}\min_{p_{1}}\chi(p_{3},\mathbf{W}(p_{3},p_{1}))$
+or
+$\max_{p_{3}}\min_{p_{2}}\chi(p_{3},\mathbf{W}(p_{3},p_{2}))\text{ .}$
+This gives us the upperbound.
 The same technique can be extended to a Byzantine multiple-access channel with $k$ senders and at most one adversarial sender:
-Corollary III.3.
+Theorem III.3.
 Let $\mathtt{W}$ be a classical-quantum $k$-user Byzantine multiple-access channel with at most one adversarial sender.
@@
 For each sender $i\in\{1,\dots,k\}$, we denote the arbitrarily varying classical-quantum channel that maps the message of sender $i$ into the output state after having already measured the messages of senders $11,\dots,i-$ 1 by $\mathtt{W}_{i}^{j}$, where $x^{n}_{j}$ and $\gamma_{j}$ represent the jamming attack by the adversarial sender $j\in\{1,\dots,i-1\}$.
-The random capacity region of $\mathtt{W}$ is included in
-$R_{i}\leq\min\Bigl{(}\max_{p_{i}}\min_{x^{n}_{j}:j<i}I(p_{i},E_{j}B_{x^{n}_{j}}|\{p_{h}:h\in\{1,\dots,i-1\}\setminus j\}),$ (5)
-$\max_{p_{i}}\min_{x^{n}_{j}:j>i}I(p_{i},B_{x^{n}_{j}}|\{p_{h}:h\in\{1,\dots,i-1\}\})\Bigr{)}\text{ .}$ (6)
+The random capacity region of $\mathtt{W}$ is
+$R_{i}\leq\min\Bigl{(}\max_{p_{i}}\min_{p_{j}:j<i}I(p_{i},E_{j}B|\{p_{h}:h\in\{1,\dots,i-1\}\setminus j\}),$ (5)
+$\max_{p_{i}}\min_{p_{j}:j>i}I(p_{i},B|\{p_{h}:h\in\{1,\dots,i-1\}\})\Bigr{)}\text{ .}$ (6)
 Proof.
@@
 $\{\mathbf{W}(\cdot,t):t\in\Theta_{i}\}:X_{i}\rightarrow X_{1}\times\cdots\times X_{i-1}\otimes\mathcal{S}(A)$
-defined by
+denoted by
 $\mathbf{W}(x^{n}_{i},x^{n}_{j})$
@@
 When
-$|M_{3}|<\min_{x^{n}_{j}:j<i}I(p_{i},E_{j}B_{x^{n}_{j}}|\{p_{h}:h\in\{1,\dots,i-1\}\setminus j\})$
+$|M_{3}|<\min_{p_{j}:j<i}I(p_{i},E_{j}B|\{p_{h}:h\in\{1,\dots,i-1\}\setminus j\})$
 and
-$|M_{3}|<\min_{x^{n}_{j}:j>i}I(p_{i},B_{x^{n}_{j}}|\{p_{h}:h\in\{1,\dots,i-1\}\}),$
+$|M_{3}|<\min_{p_{j}:j>i}I(p_{i},B|\{p_{h}:h\in\{1,\dots,i-1\}\}),$
 then by [5] and [14], there exists a reliable random code for $\{\mathbf{W}(\cdot,t):t\in\Theta_{i}\}$.
+[…]
+$D_{m_{i}}=\bigotimes_{m_{1},\cdots,m_{i-1}}D_{m_{i}}^{(m_{1},\cdots,m_{i-1})}$
+for some operators $\{D_{m_{i}}^{(m_{1},\cdots,m_{i-1})}:m_{1},\cdots,m_{i-1}\}$.
 The receiver now applies a similar strategy to the quantum multiple access channel coding in [34].
@@
 This shows that the receiver can utilize the part of the code corresponding to $m_{1},\dots,m_{i-1}$ for the channel $\{\mathbf{W}(\cdot,t):t\in\Theta_{i}\}$ as an arbitrarily varying channel code in order to decode the message of sender $i$.
+The converse can be shown in a similar as the proof of the converse of Theorem III.2.
+We consider now the case when sender i sends $x^{\gamma_{i}}_{i}(m_{i})$ through the channel.
+Let us define an arbitrarily varying classical-quantum channel
+$\{\mathsf{W}(\cdot,t):t\in\Theta_{i}\}:X_{i}\rightarrow\mathcal{S}(A)$
+denoted by
+$\mathsf{W}(x^{n}_{i},x^{n}_{j})$
+$=\sum_{m_{1},\dots,m_{k}}p(x^{n}_{1}(m_{1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{i-1}(m_{i-1}),x^{n}_{i+1}(m_{i+1}),\dots,x^{n}_{k}(m_{k})|x^{n}_{i})$
+$\quad\cdot\sqrt{D_{m_{j}}}\mathtt{W}((x^{n}_{1}(m_{1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j},x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{i-1}(m_{i-1}),x^{n}_{i},x^{n}_{i+1}(m_{i+1}),\dots,x^{n}_{k}(m_{k})))\sqrt{D_{m_{j}}}$
+for $j<i$, and
+$\mathsf{W}(x^{n}_{i},x^{n}_{j})$ $=\sum_{m_{1},\dots,m_{k}}p(x^{n}_{1}(m_{1}),\dots,x^{n}_{i-1}(m_{i-1}),x^{n}_{i+1}(m_{i+1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k})|x^{n}_{i})$
+$\quad\cdot\mathtt{W}((x^{n}_{1}(m_{1}),\dots,x^{n}_{i-1}(m_{i-1}),x^{n}_{i},x^{n}_{i+1}(m_{i+1}),\dots,x^{n}_{j-1}(m_{j-1}),x^{n}_{j},x^{n}_{j+1}(m_{j+1}),\dots,x^{n}_{k}(m_{k})))$
+for $j>i$.
+When sender j is adversarial and sends $x^{n}_{j}$ as jamming attack, the expectetd value of the channel output will be closed to the arbitrarily varying classical-quantum channel output $\mathsf{W}(x^{\gamma_{i}}_{i}(m_{i}),x^{n}_{j})$.
+By [5] and [14], the rate for the message of sender i cannot exceed
+$\min_{p_{j}}I(p_{i},E_{j}B|\{p_{h}:h\in\{1,\dots,i-1\}\setminus j\})$
+when $j$ $<i$, and
+$\min_{p_{j}}I(p_{i},B|\{p_{h}:h\in\{1,\dots,i-1\}\})$
+when $j$ $<i$. This gives us the upperbound.
 As stated in Section III-A, the channel output can be altered by the decoding operator.
@@
 If $x^{n}_{1}\rightarrow\mathtt{W}((x^{n}_{1},x^{n}_{2}))$ is not symmetrizable, and $x^{n}_{2}\rightarrow\mathtt{W}((x^{n}_{1},x^{n}_{2}))$ is not orthogonally symmetrizable, then the capacity region of $\mathtt{W}$ is given by
-$R_{1}$ $\leq\max_{p_{1}}\min_{x^{n}_{2}}I(p_{1},B_{x^{n}_{2}})$ (7)
-$R_{2}$ $\leq\max_{p_{2}}\min_{x^{n}_{1}}I(p_{2},B_{x^{n}_{1}})\text{ .}$ (8)
+$R_{1}$ $\leq\max_{p_{1}}\min_{p_{2}}I(p_{1},B)$ (7)
+$R_{2}$ $\leq\max_{p_{2}}\min_{p1}I(p_{2},B)\text{ .}$ (8)
 Proof.
@@
 Additionally, we established an achievable rate region for the $k$-user Byzantine multiple-access classical-quantum channels when $k\geq 3$.
-A general upper bound for the rate region with more than three senders remains an open problem.
-Furthermore, an important challenge lies in applying advanced classical techniques, such as adversarial sender identification from [29] and [30], to quantum channels.
+An important challenge lies in applying advanced classical techniques, such as adversarial sender identification from [29] and [30], to quantum channels.
 The main difficulty, as mentioned earlier, is that for quantum arbitrarily varying channels, only random coding techniques are currently known.
-Acknowledgments
-The authors acknowledge the support of the German Federal Ministry of Education and Research (BMBF) under several national initiatives: the 6G Communication Systems initiative through the research hub 6G-life (grant 16KISK263), the QuaPhySI (Quantum Physical Layer Service Integration) initiative (grant 16KIS2234), the “QTOK - Quantum Tokens for Secure Authentication in Theory and Practice” initiative (grant 16KISQ038), the QR.N initiative (grant 16KIS2196), the QTREX initiative (grant 16KISR038), the QUIET initiative (grant 16KISQ0170), and the QD-CAMNetz initiative (grant 16ISQ169).
+Acknowledgments.
+The authors acknowledge the support of the German Federal Ministry of Education and Research (BMBF) under several national initiatives: the 6G Communication Systems initiative through the research hub 6G-life (grant 16KISK263), the QuaPhySI (Quantum Physical Layer Service Integration) initiative (grant 16KIS2234), the "QTOK - Quantum Tokens for Secure Authentication in Theory and Practice" initiative (grant 16KISQ038), the QR.N initiative (grant 16KIS2196), the QTREX initiative (grant 16KISR038), the QUIET initiative (grant 16KISQ0170), and the QD-CAMNetz initiative (grant 16ISQ169).